OpenStudy (anonymous):

Janet is three years older than her little sister Kim. Janet's brother is eight years younger than their sister Kim. The sum of all their ages is 55 years. Can someone help me turn this into an equation?

OpenStudy (tardis):

Hi @kiyawiya! Okay so start off by making the kids names into variables. For example, Janet is X, Kim is Y, and Janet's brother is Z. The first sentence says that X is 3 more than Y. So then you can say that X = Y + 3 (or Y = X - 3). Does that help you enough to continue on your own or would you like me to guide you a little further?

OpenStudy (anonymous):

Hey @TARDIS I need a little bit more help

OpenStudy (tardis):

Let's try to make all the variables equal, okay? (P.S. I kinda messed up before lol sorry, ignore the other post) X <- This will be your base variable Y = X - 3 <- This means Kim's age is equal to Janet's age minus 3 Now for Z's substitution, it's easier to look at it in word format. If Janet is 3 years older than Kim and Kim is 8 years older than Janet's brother, then that means that Janet's brother is 11 years younger than Janet. Z = X - 11 <- Janet's brother's age is equal to Janet's age minus 11 X + Y + Z = 55 <- All their ages together equal 55 __________________ X Y = X - 3 Z = x - 11 X + Y + Z = 55 __________________ I don't know if you have to find the solution for each one but here's how, just in case. Try to make the final equation just all of one variable (X, for now) by substituting the equivalent of the other equations. X + (X - 3) + (X - 11) = 55 Now to solve for X you have to isolate X on one side X = 55 - (X - 3) - (X - 11) Take the numbers out of the parenthesis and put like terms together and then try to solve for X from there on. X = 55 + 3 + 11 - X - X X = 69 - 2X 3X = 69 X = 23 therefore Janet is 23 years old. From there, just plug in 23 into the other equations above to get Kim and Janet's brother's age.